OpenStudy (anonymous):

Turn 0.333......... Into a fraction and percent

OpenStudy (anonymous):

Percent : 33.3 %

OpenStudy (anonymous):

In fraction: 0.333 = 0.3(bar) = 3/9...

OpenStudy (anonymous):

Thanks

ganeshie8 (ganeshie8):

another way to convert .333.. to fraction : \[0.333... \\ = \frac{0.333... }{1} \\ = \frac{3*0.333.. }{3} \\ = \frac{0.999... }{3} \\ = \frac{1}{3}\]

OpenStudy (anonymous):

That's a cool stragy ! Thanks

ganeshie8 (ganeshie8):

you can find proof for 0.999... = 1 @ http://www.mathsisfun.com/9recurring.html

ganeshie8 (ganeshie8):

yw.. :))